OpenStudy (anonymous):

PLEASE HELP f(x)=17x/5lxl (absolute value of x) f(r^4)=?

OpenStudy (anonymous):

exactly what are we finding?

OpenStudy (anonymous):

f(r^4)=17(r^4)/5lr^4l (absolute value of r^4, i dont know how to type that

OpenStudy (anonymous):

Im not sure weather or not i got the right answer or not but maybe your wording is wrong cause i cant seem to find or get a answer

OpenStudy (anonymous):

hmm if i attach a word doc would you be able to open it

OpenStudy (aum):

\[ f(x) = \frac{17x}{5|x|} \\ f(r^4) = \frac{17r^4}{5|r^4|} \\ \text{r^4 is always positive or zero. It is never negative for real values of r. }\\ \text{Therefore, the absolute value can be removed. }\\ f(r^4) = \frac{17r^4}{5r^4} = \frac{17}{5}\\ \]as long as \(r \ne 0\).

OpenStudy (anonymous):

okay i see. thank you!